# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v4

@@ Abstract
[…]
@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture [52] have revolutionized natural language processing. However, their immense scale, reaching hundreds of billions or trillions of parameters  [37, 22], leads to prohibitive computational and memory costs, particularly GPU VRAM, hindering their widespread deployment on consumer or edge devices [66, 38, 63].
Model quantization, which reduces numerical precision, is a key technique to address these bottlenecks [20]. While post-training quantization (PTQ) methods like GPTQ [11] and AWQ [30] compress to $\sim$4 bits, further compression (e.g., to 1-bit) typically requires quantization-aware training (QAT) [23, 27] to maintain performance. Indeed, QAT has demonstrated the ability to enable substantial 1-bit compression while preserving model fidelity [53, 57, 32].
However, even 1-bit models (e.g., $\sim$15.4GB for 70B parameters [20]) can exceed the capacity of highly resource-constrained devices [67], motivating the exploration of sub-1-bit quantization. Although prior work [10] has ventured into this area, maintaining performance at extremely low effective bits (e.g., 0.55 bits per weight (BPW)) while ensuring computational efficiency without restrictive hardware dependencies remains a critical challenge [10]. Consequently, it is crucial to pioneer new techniques that can achieve such profound model compression, aiming for bit rates around 0.1 BPW, while maintaining the integrity of model performance.
@@
Our approach stems from two observations. First, LLM weight matrices often exhibit significant low-rank structure [24, 54]. This suggests that factorization methods like Singular Value Decomposition (SVD) [14] could offer a more stable compression pathway than pruning [17], especially under extreme compression regimes. For instance, at compression ratios exceeding 50%, SVD-based approaches [55] consistently outperform pruning methods like Wanda [48] (see Section˜A.5 for details). Second, binarization inherently causes information loss [15]. Recent high-performing 1-bit methods [57, 25, 29] demonstrate that sophisticated scaling over multiple dimensions (e.g., row and column), is crucial for mitigating this loss and stabilizing performance [29].
Based on these insights, we introduce LittleBit, a novel method for extreme sub-1-bit quantization (e.g., 0.1 BPW). LittleBit first represents weights via low-rank factorization ($\mathbf{W}\approx\mathbf{U}\mathbf{V}^{\top}$) and then binarizes these factors. To counteract errors from binarization, LittleBit leverages a multi-scale compensation mechanism that applies learnable scales across rows, columns, and an additional latent dimension. This entire architecture is then complemented by Residual Compensation.
Extensive experiments show LittleBit’s superiority over STBLLM [10], the leading sub-1-bit technique, on LLMs from 1.3B to 32B parameters. Notably, LittleBit achieves competitive performance on Llama2-13B at an unprecedented 0.1 BPW (Fig.˜1), surpassing STBLLM at 0.55 BPW. Furthermore, the 32B LittleBit model at just 0.3 BPW maintains strong performance, significantly outperforming STBLLM and setting a new state-of-the-art for sub-1-bit quantization.
In summary, our contribution is as follows:
  • We propose LittleBit, a novel method unifying latent matrix factorization with multi-scale compensation for extreme sub-1-bit quantization. Our method achieves unprecedentedly low effective bits (e.g., down to 0.1 BPW) while preserving model performance, establishing a new, viable size-performance trade-off that provides a practical path for deploying LLMs in highly resource-constrained environments.
  • We introduce Dual-SVID for effective initialization of factorized structures and Residual Compensation to mitigate errors in extreme quantization.
  • We provide extensive validation showing that LittleBit significantly outperforms STBLLM [10], the prior leading sub-1-bit quantization technique, on various LLM scales.
@@ 2.1 Binarization and Quantization
Network binarization targets extreme compression and acceleration by constraining weights or activations to $\pm 1$, enabling efficient bitwise operations [21, 12]. Although early works like BinaryConnect [7] and BNNs [21] showed feasibility, they faced significant accuracy degradation [39]. This degradation was partially mitigated by introducing scaling factors [41] and employing the Straight-Through Estimator (STE) [3] to handle the non-differentiable quantization function during training [44].
However, directly applying these early techniques to sensitive large language models (LLMs) typically results in unacceptable performance loss [45]. To address this challenge, LLM-specific quantization strategies have been developed, mainly categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [13]. PTQ adapts pre-trained models, achieving competitive results around 4-bit precision with methods such as GPTQ [11] and AWQ [30], but generally struggles to maintain performance below 2-bit precision [45, 20]. The sub-1-bit regime is particularly challenging for PTQ methods due to severe information loss [9]. Recent work such as STBLLM [10] shows that incorporating structured binarized compression can push PTQ into the sub-1-bit regime, yet a non-negligible accuracy gap remains.
In contrast, QAT integrates the quantization process directly into the training loop. This enables the model to learn and adapt to the low-precision format during training, generally yielding superior performance compared to PTQ at very low bit-widths [31]. Successful QAT approaches for low-bit scenarios often employ sophisticated scaling mechanisms alongside the adaptive training itself, as evidenced by several recent methods [57, 25]. However, despite these advancements, consistently achieving high model fidelity when quantizing to less than 1.0 BPW remains a considerable challenge. This persisting difficulty underscores the need for novel architectures and training strategies specifically designed for such extreme compression regimes.
@@ 2.2 Low-Rank Approximation in Quantization
Beyond quantization, the inherent parameter redundancy within deep neural networks [9], particularly evident in LLMs [34, 18], can be effectively exploited using low-rank approximation methods. Techniques such as Singular Value Decomposition (SVD) [14] allow large weight matrices to be represented as products of smaller factors, offering an alternative compression strategy that can be more resilient than pruning, especially under high compression regimes [17]. Recognizing their potential for complementary benefits, researchers have explored combining low-rank factorization to reduce parameter count and quantization to lower parameter precision [42]. Initial studies applied SVD in various roles within the LLM context: some used activation statistics to guide decomposition [55], others integrated SVD principles into initialization or parameter-efficient updates [61, 62, 57], applied transformations prior to factorization [58], or even employed low-rank structures to model quantization errors [60].
@@
A more integrated approach involves incorporating low-rank concepts directly within the QAT methods, as demonstrated by DL-QAT [26], which jointly learns a representation amenable to both low-rank approximation and low-precision quantization, albeit limited to 3-bit quantization. This direction of optimizing structural properties such as rank and numerical precision during training represents a promising path towards more effective LLM compression, with potential for further exploration of even lower precision quantization.
@@ 3 Methodology
This section details LittleBit. We describe its factorized architecture with multi-scale compensation (Section˜3.1), the Dual-SVID initialization for stable QAT (Section˜3.2), and Residual Compensation for enhanced fidelity (Section˜3.3). These components are optimized via QAT (details in Section˜4).
@@ 3.1 LittleBit Architecture
The LittleBit architecture, depicted in Fig.˜2, redesigns linear layers for extreme compression (e.g., 0.1 BPW) by synergistically combining low-rank factorization and binarization. LittleBit first leverages the observed low-rank structure in LLM weights [56] to approximate the weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ as:
@@
This factorization-based approach is chosen over other structural compression methods, such as pruning, due to its superior robustness in extreme compression regimes. As demonstrated in our comparative analysis against methods like Wanda (Section˜A.5), SVD-based compression maintains model integrity far more effectively at high compression ratios, providing a more stable foundation for our ultra low-bit quantization. The resulting factors $\mathbf{U}$ and $\mathbf{V}$ are then binarized:
@@
To compensate for the significant information loss incurred by binarization [15], LittleBit incorporates learnable FP16 scales. Beyond standard row ($\mathbf{h}$) and column ($\mathbf{g}$) scaling [45, 57, 29], it introduces an additional latent scale ($\bm{\ell}\in\mathbb{R}^{r}$). This latent scale is designed to address the factorization structure by learning the relative importance of each of the $r$ latent dimensions, which correspond to the columns of the binarized factors $\mathbf{U}_{\mathrm{sign}}$ and $\mathbf{V}_{\mathrm{sign}}$:
@@
Instead of storing or optimizing $\widehat{\mathbf{W}}_{\mathrm{pri}}$ directly, LittleBit learns latent full-precision factors $\mathbf{U}$ and $\mathbf{V}$, which are binarized during the forward pass using Eq.˜2, along with the FP16 scales $\mathbf{h},\mathbf{g},\bm{\ell}$. This factorized representation enables efficient computation during the forward pass.
@@
Let $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$ be the input matrix. The forward computation $\mathbf{Y}=\mathbf{X}{\widehat{\mathbf{W}}_{\mathrm{pri}}}^{\top}$ using the primary approximation (Eq.˜4) can be efficiently computed as:
@@
This decomposition (Eq.˜5) replaces a large high precision General Matrix Multiply (GEMM) operation with two smaller binary matrix multiplications and element-wise scaling operations, offering computational advantages and reducing memory requirements. The effective bits per weight (BPW) is determined by the storage cost of these learnable parameters relative to the original matrix size (see Section˜A.4 for more details).
@@ 3.2 Dual-SVID Initialization
Naively initialing the highly constrained LittleBit structure can lead to unstable QAT. To circumvent this, we propose Dual-SVID, an SVD-based initialization method designed to provide a starting point where the initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ closely approximates $\mathbf{W}$. Dual-SVID aims to preserve essential sign and magnitude information from the optimal low-rank SVD factors (obtained from $\mathbf{W}\approx\mathbf{U}^{\prime}{\mathbf{V}^{\prime}}^{\top}$, then truncated) and map this information to the learnable LittleBit parameters ($\mathbf{U}_{\mathrm{sign}},\mathbf{V}_{\mathrm{sign}},\mathbf{h},\mathbf{g},\bm{\ell}$). The process involves: (1) Obtaining $\mathbf{U}^{\prime},\mathbf{V}^{\prime}$ via a truncated SVD. (2) Setting initial binary factors based on signs (Eq.˜6). (3) Decomposing the magnitudes $|\mathbf{U}^{\prime}|\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $|\mathbf{V}^{\prime}|\in\mathbb{R}^{d_{\mathrm{in}}\times r}$. To initialize the scales, we perform a rank-1 SVD (or equivalent rank-1 approximation) on each of these magnitude matrices separately. For $|\mathbf{U}^{\prime}|$, its best rank-1 approximation can be written as $\mathbf{s}_{U,0}(\bm{\ell}_{u,0})^{\top}$, where $\mathbf{s}_{U,0}\in\mathbb{R}^{d_{\mathrm{out}}}$ becomes the initial row scale $\mathbf{h}_{0}$, and $\bm{\ell}_{u,0}\in\mathbb{R}^{r}$ is a component contributing to the latent scale. Similarly, for $|\mathbf{V}^{\prime}|$, its rank-1 approximation $\mathbf{s}_{V,0}(\bm{\ell}_{v,0})^{\top}$ provides the initial column scale $\mathbf{g}_{0}=\mathbf{s}_{V,0}\in\mathbb{R}^{d_{\mathrm{in}}}$ and another latent scale component $\bm{\ell}_{v,0}\in\mathbb{R}^{r}$. These steps effectively disentangle multi-dimensional magnitude variations and allow us to estimate initial row ($\mathbf{h}_{0}$), column ($\mathbf{g}_{0}$), and the final latent scale ($\bm{\ell}_{0}=\bm{\ell}_{u,0}\odot\bm{\ell}_{v,0}$ by element-wise product), as shown in (Eqs.˜7 and 8):

### v5

@@ Abstract
The deployment of large language models (LLMs) is frequently hindered by prohibitive memory and computational requirements. While quantization mitigates these bottlenecks, maintaining model fidelity in the sub-1-bit regime remains a persistent challenge. In this paper, we introduce LittleBit, a novel framework for extreme LLM compression. We target quantization rates as low as $0.1$ bits per weight (BPW), achieving a memory reduction of approximately $31\times$, which effectively compresses Llama2-13B to under $0.9$ GB. We represent weights via low-rank latent matrix factorization and subsequently binarize the resulting factors. To counteract the information loss inherent to such drastic precision reduction, we integrate a multi-scale compensation mechanism that learns importance parameters across row, column, and latent dimensions. Two primary contributions enable effective training: Dual Sign-Value-Independent Decomposition (Dual-SVID) for quantization-aware training (QAT) initialization, and Residual Compensation to minimize approximation errors. Extensive experiments confirm the superiority of LittleBit in the sub-1-bit domain; for instance, our method at $0.1$ BPW surpasses the performance of leading techniques operating at $0.7$ BPW on Llama2-7B. We establish a new size-performance trade-off—unlocking a potential $11.6\times$ inference speedup relative to FP16—and render powerful LLMs practical for resource-constrained environments. Our code is available at https://github.com/SamsungLabs/LittleBit.
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture [53] have fundamentally transformed natural language processing. However, the scale of these models, often reaching hundreds of billions or trillions of parameters [38, 22], results in prohibitive computational and memory costs. In particular, the high demand for GPU VRAM hinders widespread deployment on consumer or edge devices [66, 39, 62].
Model quantization, which reduces numerical precision, serves as a primary technique to address these bottlenecks [20]. While post-training quantization (PTQ) methods such as GPTQ [11] and AWQ [30] effectively compress models to approximately 4 bits, achieving further compression (e.g., to 1-bit) typically requires quantization-aware training (QAT) [23, 27] to maintain performance. QAT has demonstrated the capability to enable 1-bit compression while preserving model fidelity [54, 58, 32].
Nevertheless, even 1-bit models (e.g., approximately 15.4 GB for 70B parameters [20]) may exceed the capacity of highly resource-constrained devices [67]. This limitation motivates the exploration of sub-1-bit quantization. Although prior work [10] has investigated this area, maintaining performance at extremely low effective bits (e.g., 0.55 bits per weight (BPW)) while ensuring computational efficiency without restrictive hardware dependencies remains challenging [10]. Consequently, it is necessary to develop new techniques that achieve such extreme model compression, aiming for bit rates around 0.1 BPW, while preserving model performance.
@@
We derive our approach from two observations. First, LLM weight matrices often exhibit pronounced low-rank structure [24, 55]. This observation suggests that factorization methods such as Singular Value Decomposition (SVD) [14] may offer a more stable compression pathway than pruning [17], particularly under extreme compression regimes. For instance, at compression ratios exceeding 50%, SVD-based approaches [56] consistently outperform pruning methods such as Wanda [49] (see Appendix˜E for details). Second, binarization inherently causes information loss [15]. Recent high-performing 1-bit methods [58, 25, 29] demonstrate that comprehensive scaling over multiple dimensions (e.g., row and column) is crucial for mitigating this loss and stabilizing performance [29].
Building on these insights, we introduce LittleBit, a novel method for extreme sub-1-bit quantization (e.g., 0.1 BPW). In LittleBit, we first represent weights via low-rank factorization ($\mathbf{W}\approx\mathbf{U}\mathbf{V}^{\top}$) and subsequently binarize these factors. To counteract errors from binarization, we leverage a multi-scale compensation mechanism that applies learnable scales across rows, columns, and an additional latent dimension. We complement this architecture with Residual Compensation.
Extensive experiments demonstrate the superiority of LittleBit over STBLLM [10], the leading sub-1-bit technique, on LLMs ranging from 1.3B to 32B parameters. Notably, LittleBit achieves competitive performance on Llama2-13B at 0.1 BPW (Fig.˜1), surpassing STBLLM at 0.55 BPW. Furthermore, the 32B LittleBit model at just 0.3 BPW maintains strong performance, substantially outperforming STBLLM and establishing a new state-of-the-art for sub-1-bit quantization.
In summary, our contributions are as follows:
  • We propose LittleBit, a novel framework unifying latent matrix factorization with multi-scale compensation to enable extreme sub-1-bit quantization. This method achieves effective bits down to 0.1 BPW while preserving model performance, establishing a viable size-performance trade-off for deploying LLMs in resource-constrained environments.
  • We introduce Dual-SVID for the effective initialization of factorized structures and integrate Residual Compensation to mitigate approximation errors. These techniques collectively stabilize quantization-aware training (QAT) and counteract the information loss inherent in drastic precision reduction, ensuring robust learning dynamics.
  • We provide extensive empirical validation demonstrating that LittleBit consistently outperforms STBLLM [10], the prior leading sub-1-bit technique, across various model scales. Our results confirm that LittleBit maintains superior fidelity in the deep compression regime, where baseline methods typically suffer severe degradation.
@@ 2.1 Binarization and Quantization
Network binarization aims for extreme compression and acceleration by constraining weights or activations to $\pm 1$. This enables efficient bitwise operations [21, 12]. While early studies such as BinaryConnect [7] and BNNs [21] demonstrated feasibility, they encountered substantial accuracy degradation [40]. Researchers partially mitigated this degradation by introducing scaling factors [42] and employing the Straight-Through Estimator (STE) [3] to handle the non-differentiable quantization function during training [45].
The direct application of these early techniques to large language models (LLMs) typically results in severe performance loss [46]. Consequently, LLM-specific quantization strategies have emerged. These are primarily categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [13]. PTQ adapts pre-trained models and achieves competitive results at approximately 4-bit precision with methods such as GPTQ [11] and AWQ [30]. However, these methods generally struggle to maintain performance below 2-bit precision [46, 20]. The sub-1-bit regime presents particular challenges for PTQ methods due to severe information loss [9]. Recent work such as STBLLM [10] demonstrates that incorporating structured binarized compression can extend PTQ into the sub-1-bit regime, but a non-negligible accuracy gap persists.
In contrast, QAT integrates the quantization process directly into the training loop. This enables the model to learn and adapt to the low-precision format [31]. This approach generally yields superior performance compared to PTQ at very low bit-widths. Successful QAT approaches for low-bit scenarios often employ sophisticated scaling mechanisms alongside adaptive training, as evidenced by recent methods [58, 25]. Despite these advancements, consistently achieving high model fidelity when quantizing to less than $1.0$ BPW remains a considerable challenge. This persisting difficulty underscores the need for novel architectures and training strategies specifically designed for such extreme compression regimes.
@@ 2.2 Low-Rank Approximation in Quantization
Beyond quantization, the inherent parameter redundancy within deep neural networks [9] can be effectively exploited using low-rank approximation methods. This redundancy is particularly evident in LLMs [34, 18]. Techniques such as Singular Value Decomposition (SVD) [14] allow large weight matrices to be represented as products of smaller factors. This offers an alternative compression strategy that may be more resilient than pruning, especially under high compression regimes [17]. Recognizing the potential for complementary benefits, researchers have explored combining low-rank factorization to reduce parameter count with quantization to lower parameter precision [43]. Initial studies applied SVD in various roles within the LLM context. Some approaches used activation statistics to guide decomposition [56], while others integrated SVD principles into initialization or parameter-efficient updates [63, 35, 58]. Further works applied transformations prior to factorization [59] or employed low-rank structures to model quantization errors [61].
@@
A more integrated approach involves incorporating low-rank concepts directly within QAT methods. For example, DL-QAT [26] jointly learns a representation amenable to both low-rank approximation and low-precision quantization. However, this method is currently limited to 3-bit quantization. Optimizing structural properties such as rank and numerical precision during training represents a promising path toward effective LLM compression. There is potential for further exploration of even lower precision quantization.
@@ 3 Methodology
This section details LittleBit. It describes the factorized architecture with multi-scale compensation (Section˜3.1), the Dual-SVID initialization for stable QAT (Section˜3.2), and Residual Compensation for enhanced fidelity (Section˜3.3). These components are optimized via QAT, as detailed in Section˜4.
@@ 3.1 LittleBit Architecture
The LittleBit architecture, depicted in Fig.˜2, redesigns linear layers for extreme compression (e.g., 0.1 BPW) by synergistically combining low-rank factorization and binarization. The method leverages the observed low-rank structure in LLM weights [57] to approximate the weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ as:
@@
We prioritize this factorization-based approach over other structural compression methods, such as pruning, due to its superior robustness in extreme compression regimes. As the comparative analysis against methods like Wanda demonstrates (Appendix˜E), SVD-based compression maintains model integrity far more effectively at high compression ratios. This provides a more stable foundation for ultra low-bit quantization. The resulting factors $\mathbf{U}$ and $\mathbf{V}$ are then binarized:
@@
To compensate for the significant information loss incurred by binarization [15], LittleBit incorporates learnable FP16 scales. Beyond standard row ($\mathbf{h}$) and column ($\mathbf{g}$) scaling [46, 58, 29], the architecture introduces an additional latent scale ($\bm{\ell}\in\mathbb{R}^{r}$). This latent scale addresses the factorization structure by learning the relative importance of each of the $r$ latent dimensions, which correspond to the columns of the binarized factors $\mathbf{U}_{\mathrm{sign}}$ and $\mathbf{V}_{\mathrm{sign}}$:
@@
Instead of storing or optimizing $\widehat{\mathbf{W}}_{\mathrm{pri}}$ directly, the model learns latent full-precision factors $\mathbf{U}$ and $\mathbf{V}$. These are binarized during the forward pass using Eq.˜2, along with the FP16 scales $\mathbf{h},\mathbf{g},\bm{\ell}$. This factorized representation enables efficient computation during the forward pass.
@@
Let $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$ be the input matrix. The forward computation $\mathbf{Y}=\mathbf{X}{\widehat{\mathbf{W}}_{\mathrm{pri}}}^{\top}$ using the primary approximation (Eq.˜4) is efficiently computed as:
@@
This decomposition (Eq.˜5) replaces a large high-precision General Matrix Multiply (GEMM) operation with two smaller binary matrix multiplications and element-wise scaling operations. This offers computational advantages and reduces memory requirements. The effective bits per weight (BPW) is determined by the storage cost of these learnable parameters relative to the original matrix size (see Appendix˜D for details).
@@ 3.2 Dual-SVID Initialization
Naively initializing the highly constrained LittleBit structure can lead to unstable QAT. To circumvent this, we propose Dual-SVID, an SVD-based initialization method designed to provide a starting point where the initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}$ closely approximates $\mathbf{W}$. Dual-SVID aims to preserve essential sign and magnitude information from the optimal low-rank SVD factors (obtained from $\mathbf{W}\approx\mathbf{U}^{\prime}{\mathbf{V}^{\prime}}^{\top}$, then truncated) and map this information to the learnable LittleBit parameters ($\mathbf{U}_{\mathrm{sign}},\mathbf{V}_{\mathrm{sign}},\mathbf{h},\mathbf{g},\bm{\ell}$). The process involves three steps: (1) Obtaining $\mathbf{U}^{\prime},\mathbf{V}^{\prime}$ via a truncated SVD. (2) Setting initial binary factors based on signs (Eq.˜6). (3) Decomposing the magnitudes $|\mathbf{U}^{\prime}|\in\mathbb{R}^{d_{\mathrm{out}}\times r}$ and $|\mathbf{V}^{\prime}|\in\mathbb{R}^{d_{\mathrm{in}}\times r}$. To initialize the scales, a rank-1 SVD (or equivalent rank-1 approximation) is performed on each of these magnitude matrices separately. For $|\mathbf{U}^{\prime}|$, its best rank-1 approximation can be written as $\mathbf{s}_{U,0}(\bm{\ell}_{u,0})^{\top}$, where $\mathbf{s}_{U,0}\in\mathbb{R}^{d_{\mathrm{out}}}$ becomes the initial row scale $\mathbf{h}_{0}$, and $\bm{\ell}_{u,0}\in\mathbb{R}^{r}$ is a component contributing to the latent scale. Similarly, for $|\mathbf{V}^{\prime}|$, its rank-1 approximation $\mathbf{s}_{V,0}(\bm{\ell}_{v,0})^{\top}$ provides the initial column scale $\mathbf{g}_{0}=\mathbf{s}_{V,0}\in\mathbb{R}^{d_{\mathrm{in}}}$ and another latent scale component $\bm{\ell}_{v,0}\in\mathbb{R}^{r}$. These steps effectively disentangle multi-dimensional magnitude variations and allow for the estimation of initial row ($\mathbf{h}_{0}$), column ($\mathbf{g}_{0}$), and the final latent scale ($\bm{\ell}_{0}=\bm{\ell}_{u,0}\odot\bm{\ell}_{v,0}$ by element-wise product), as shown in Eqs.˜7 and 8:
@@
The resulting initial effective weight $\widehat{\mathbf{W}}_{\mathrm{pri},0}=\mathrm{diag}(\mathbf{h}_{0})\mathbf{U}_{\mathrm{sign},0}\mathrm{diag}(\bm{\ell}_{0})\mathbf{V}_{\mathrm{sign},0}^{\top}\mathrm{diag}(\mathbf{g}_{0})$ preserves key structural information. The learnable parameters are initialized as follows: the latent factors $\mathbf{U}$ and $\mathbf{V}$ are initialized with the factors $\mathbf{U}^{\prime}$ and $\mathbf{V}^{\prime}$ obtained from SVD, respectively, and the scales $\mathbf{h},\mathbf{g},\bm{\ell}$ are initialized to $\mathbf{h}_{0},\mathbf{g}_{0},\bm{\ell}_{0}$.
3.3 Residual Compensation
To further improve fidelity without increasing the primary path’s rank, we introduce Residual Compensation. This technique is motivated by the theoretical insight that a two-stage error correction can be more effective than a single, larger approximation (Appendix˜A). Crucially, this method does not increase the model’s total parameter budget; instead, the fixed bit budget of a single, higher-rank approximation is strategically reallocated into two lower-rank paths: a primary and a residual path. The term ‘residual’ primarily describes its role during initialization, where the auxiliary path is configured to model the error of the primary approximation. During QAT, both paths are optimized jointly to collectively represent the original weight. The advantage of this dual-path approach over a single path is empirically validated in our ablation studies (Section˜B.1).